Find the electrostatic force between two point charges of equal magnitude

12μC each separated by 10cm of distance.


1
Expert's answer
2021-04-16T07:22:07-0400
"F=k\frac{q^2}{r^2}=9·10^9·\frac{(12·10^{-6})^2}{0.1^2}=129.4\text{ N}."